A company produces the financial results shown in the table below. The executives at the firm have good reason to believe that $10.3 million in sales will be generated in 2010. Using a simple linear regression model, forecast the amount of profit they might expect in 2010. Be sure to compute the associated R-squared value and provide a related managerial interpretation of it.

Year

Sales Totals (in millions)

Profit Totals (in millions)

1998

$7.0

$0.15

1999

$2.0

$0.10

2000

$6.0

$0.13

2001

$4.0

$0.15

2002

$14.0

$0.25

2003

$15.0

$0.27

2004

$16.0

$0.24

2005

$12.0

$0.20

2006

$14.0

$0.27

2007

$20.0

$0.44

2008

$15.0

$0.34

2009

$7.0

$0.17
Answer #1

Use the following computation to get the estimates for slope, intercept, and the R-squared value.

So,

The regression equation will be written as follows:

Profit = 0.05060 + 0.01593 * Sales

So, if Sales = 10.3 million, Profit = 0.05060 + 0.01593*10.3 = $0.215 million

The R-squared value of 0.8403. This means that 84.03% of the total variations of 'Profit' can be explained by 'Sales' using the above model.